OpenStudy (anonymous):

There are five independent stages in the repair of a certain piece of computer equipment.The repair time for each stage is exponentially distributed with average value of 10minutes.What is the probability that a customer engineer can repair the equipment in an hour or less?Not more that 90 minutes?

OpenStudy (king):

5/9??

OpenStudy (anonymous):

Can you tell me how did u get that ?

OpenStudy (king):

Is it rite?

OpenStudy (anonymous):

Ive got a diff answer, ...which method did u use?

OpenStudy (king):

what answer u got and by which method?I got 2 answers by using 2 diff. methods..

OpenStudy (anonymous):

As it said "exponentially distributed" , i've used the exponential distribution \[1-e ^{-\lambda x}\]

OpenStudy (anonymous):

So how did u solve it?

OpenStudy (king):

i used the exponential distribution method for 1 answer and for the other i used the normall probability method by finding out sample space,no.of jobs etc....

OpenStudy (anonymous):

Hmmm...but i guess its only the expo. dist. that we need to use...the 1st bit askes us to find P{X<=60} and 2nd one P{X<90}

OpenStudy (anonymous):

For F(x)=\[1-e ^{-\lambda x}\] , where i got lambda=0.1 ( from given avg=10)

OpenStudy (king):

0.10?

OpenStudy (king):

1/10?

OpenStudy (anonymous):

yeah, since in expo. dist. mean=1/\[\lambda\] given mean=10 (avg=mean) and hence \[\lambda\] =1/10=0.1

OpenStudy (anonymous):

Okay, so what was ur answer?? for 1st bit i've got the answer as 0.99752 , which is the prob that a stage can be reapired, but the question asks "complete equipment" and that s where i got stuck...

OpenStudy (anonymous):

*repaired

OpenStudy (king):

4.287...but i shall check with my sir and let u know....

OpenStudy (anonymous):

How is it? did u do this: P{X<=60}= 1- e\[^{-(0.1)(60)}\] = 0.99752 can you plz type how have you got 4.28?? coz if u know, probability always lies between 0 and 1

OpenStudy (king):

i know that's why i wanted to ask my sir...

OpenStudy (anonymous):

Okay....but u need the answer today...gotta submit tomorrow

OpenStudy (anonymous):

*I need it